# Generating Diverse and High-Quality Texts by Minimum Bayes Risk Decoding

### v1

@@ Abstract
One of the most important challenges in text generation systems is to produce outputs that are not only correct but also diverse. Recently, Minimum Bayes-Risk (MBR) decoding has gained prominence for generating sentences of the highest quality among the decoding algorithms. However, existing algorithms proposed for generating diverse outputs are predominantly based on beam search or random sampling, thus their output quality is capped by these underlying methods. In this paper, we investigate an alternative approach – we develop diversity-promoting decoding algorithms by enforcing diversity objectives to MBR decoding. We propose two variants of MBR, Diverse MBR (DMBR) and $k$-medoids MBR (KMBR), methods to generate a set of sentences with high quality and diversity. We evaluate DMBR and KMBR on a variety of directed text generation tasks using encoder-decoder models and a large language model with prompting. The experimental results show that the proposed method achieves a better trade-off than the diverse beam search and sampling algorithms overall.
@@ 1 Introduction
In this paper, we develop a diverse decoding method by extending Minimum Bayes Risk (MBR) decoding Goel and Byrne (2000); Kumar and Byrne (2002, 2004); Eikema and Aziz (2022). MBR decoding is shown to generate higher quality sentences than random sampling and beam search in directed text generation tasks including machine translation, text summarization, and data-to-text Freitag et al. (2022); Suzgun et al. (2023). The procedure of the MBR decoding consists of the following steps. First, it samples a set of candidate outputs from the probability model. Then, it computes the similarity of each sequence to the others according to a utility function. Finally, it selects the sequence that maximizes the expected utility over the sequences. A naive approach to generate $k$ outputs using MBR is to select the top-$k$ outputs with the highest expected utility. However, it tends to select a set of similar sentences with a large overlap (see Appendix B for examples).
@@ 2 Background
In this paper, we consider a variant of the set decoding problem Meister et al. (2021b) where the task is to generate a set of $k$ sentences $H$ that maximizes the sum of $P_{\mathrm{human}}$ and the diversity according to the preference of human $d_{\mathrm{human}}$:
@@
In this paper, we use automated evaluation metrics to approximate $d_{\mathrm{human}}$. In particular, we consider pairwise-BLEU, distinct-n, and pairwise-sentence BERT as measures of diversity Shen et al. (2019); Li et al. (2016); Reimers and Gurevych (2019).
@@ 2.1 Decoding Algorithms for Diversity
Random Sampling
Random sampling is commonly used to generate diverse outputs in both directed and open-ended text generation tasks. A simple solution is to use an ancestral sampling with a temperature parameter to control the stochasticity of the sampling. There are a lot of studies on biasing the ancestral sampling to generate higher quality outputs while maintaining the diversity of the randomized algorithm. Prior work shows that top-$k$ sampling that restricts the sampling to the $k$ most likely tokens at each step is a better alternative to controlling the temperature in a story generation task Fan et al. (2018); Ippolito et al. (2019). Nucleus sampling Holtzman et al. (2020) is similar to top-$k$ sampling and has shown to be more effective than top-$k$ sampling in WebText dataset Radford et al. (2019). Nucleus sampling truncates all tokens except those in the nucleus, the smallest possible set of tokens that covers a fraction $p$ of the model probability. Epsilon sampling Hewitt et al. (2022) is also a variant of ancestral sampling that truncates tokens whose probability is less than a threshold. They are shown to be more effective than nucleus sampling in the WebText dataset.
Variants of Beam Search
Another line of work is to generate a diverse set of outputs by promoting diversity objectives to the beam search. Beam search considers a fixed number of $k$ candidate sequences at each step Graves (2012); Sutskever et al. (2014). It is known to produce higher quality sequences than random sampling in a wide range of tasks. Diversity is induced in the beam search procedure in various forms. Li and Jurafsky (2016) propose to add a diversity constraint to standard beam search so that the number of descendants from the same parent hypothesis is capped at some upper bound. Noisy Parallel Approximate Decoding induces noise to the hidden state of the decoder to generate randomized outputs Cho (2016). Diverse beam search (DBS) adds a diversity term to the reranking objective, penalizing sequences with a small Hamming distance to other groups of the sequences Vijayakumar et al. (2016, 2018). Iterative beam search runs beam search multiple times with a constraint that any partial hypothesis that has been generated in previous iterations is prohibited Kulikov et al. (2019). Clustered beam search prunes similar sequences by clustering the candidates using a word embedding at each decoding step Tam (2020). Post-Decoding Clustering (PDC) proposes to encourage diversity by running the clustering after generating a large number of outputs, selecting good candidates from each cluster Kriz et al. (2019); Ippolito et al. (2019). Best-$k$ search points out that the greedy nature of the beam search is the source of the low diversity and shows that the best-first policy can maintain a diverse set of hypotheses as candidates, resulting in higher quality and diversity outputs Xu et al. (2023).
@@ 2.2 Minimum Bayes Risk (MBR) Decoding
One of the most common decision rules to solve the decoding problem is maximum-a-posteriori (MAP) decoding. MAP decoding finds the most probable translation under the model.
@@
$\mathbf{h}^{\mathrm{MBR}}=\operatorname*{arg\,max}_{\mathbf{h}\in\mathcal{H}_{\mathrm{cand}}}\frac{1}{N}\sum_{\mathbf{y}\in\mathcal{H}_{\mathrm{ref}}}u(\mathbf{h},\mathbf{y})$ (7)
where $f$ is the objective function of the MBR decoding and $N=|\mathcal{H}_{\mathrm{ref}}|$. Standard practice is to use the same set of hypotheses for the candidate pool ($\mathcal{H}$) and the reference pool ($\mathcal{H}_{\mathrm{ref}}$), therefore $\mathcal{H}=\mathcal{H}_{\mathrm{ref}}$.
@@ 3 Minimum Bayes Risk Decoding with Diversity
However, it results in a set of similar sentences with a large overlap. In fact, it often finds almost duplicated sentences in machine translation tasks (Table 1 and 2 in Appendix B).
@@ 3.1 Diverse MBR (DMBR)
$H^{\mathrm{DMBR}}=$ (9)
$\;\;\operatorname*{arg\,max}_{\stackrel{{\scriptstyle H\subseteq\mathcal{H}_{\mathrm{cand}}}}{{|H|=k}}}\sum_{\mathbf{h}\in H}\left(\frac{1}{N}\sum_{\mathbf{y}\in\mathcal{H}_{\mathrm{ref}}}u(\mathbf{h},\mathbf{y})\right)+d(H)$
@@
$d(H)$ is the diversity objective that penalizes according to a user-defined diversity objective. In this paper, we aim to minimize the utility between each pair of outputs so that the pairwise similarity is minimized:
@@ 3.2 $k$-Medoids MBR (KMBR)
As an alternative approach to promote diversity, we propose $k$-Medoids MBR (KMBR) decoding. $k$-Medoids is a clustering problem similar to $k$-means in that it chooses centers to minimize the total distance of data points to the closest center points. The difference is that $k$-Medoids needs to choose actual data points as centers Rdusseeun and Kaufman (1987); Kaufman and Rousseeuw (2009). Intuitively, $k$-Medoids center points are supposed to be representative of different clusters of hypotheses. Thus, picking the center points for clusters is likely to result in a set of diverse and high-quality hypotheses. KMBR can be understood as a generalization of the vanilla MBR decoding which is solving the $1$-Medoid problem to find the single most centered hypothesis out of the sampled hypotheses Jinnai and Ariu (2024). We use the negative utility as a distance and consider the problem of picking a set of hypotheses as center points. The total distance from the reference set (= data points) to the picked center points is minimized:
$H^{\mathrm{KMBR}}$ $=\operatorname*{arg\,max}_{\stackrel{{\scriptstyle H\subseteq\mathcal{H}_{\mathrm{cand}}}}{{|H|=k}}}\sum_{\mathbf{y}\in\mathcal{H}_{\mathrm{ref}}}\min_{\mathbf{h}\in H}-u(\mathbf{h},\mathbf{y})$ (11)
Note that KMBR with $k=1$ recovers the vanilla MBR decoding with a single output (Eq. 6). $k$-medoids can be used with arbitrary dissimilarity measures. Therefore, there is no restriction on the choice of the utility function $u$. Because the $k$-medoids problem is NP-hard to solve exactly, we deploy Partition Around Medoids (PAM) to compute $H^{\mathrm{KMBR}}$ approximately Park and Jun (2009). Similarly to DMBR, KMBR needs to compute the utility function for every pair of hypotheses so it is still slower than MBR even with the approximation algorithm.
@@
(b) distinct-2 $\uparrow$ (De-En)
(c) distinct-3 $\uparrow$ (De-En)
(d) P-BLEU $\downarrow$ (Ru-En)
(e) distinct-2 $\uparrow$ (Ru-En)
(f) distinct-3 $\uparrow$ (Ru-En)
Figure 1: Evaluation of pairwise-BLEU (P-BLEU) and distinct-2, 3 as a function of mean BLEU on WMT’19 De-En and Ru-En. The size of the output $k$ is set to 4. The results of distinct-1 and $k=8,12$ are present in the Appendix.
@@ 4 Experiments
We evaluate DMBR and KMBR in machine translation, text summarization, and image captioning. All experiments use BERTScore Zhang* et al. (2020) as the utility function of the MBR. We compare the performance of the sampling algorithms, beam search, diverse beam search (DBS), MBR, diverse MBR (DMBR), and $k$-Medoids MBR (KMBR).
We evaluate the quality and the diversity of the generated texts. As a quality metric, we report the mean, max, and minimum quality over each set of $k$ sentences measured by BLEU, ROUGE-L, or METEOR Papineni et al. (2002); Lin and Och (2004); Banerjee and Lavie (2005). We use a distinct-n Li et al. (2016) and a pairwise-BLEU Shen et al. (2019) as a diversity metric.
Due to limitations in computational resources, we evaluate the first 1000 entries of each dataset. We use Huggingface’s Transformers library for running all the experiments Wolf et al. (2020). We implement PAM for KMBR using the scikit-learn-extra library. [^2] https://github.com/scikit-learn-contrib/scikit-learn-extra/tree/v0.3.0 We initialize the clusters using k-medoids++ with maximum iterations set to $300$. For reproducibility, all the experiments are conducted using publicly available pretrained models and datasets. We use sacreBLEU system Post (2018) to compute BLEU scores.
(a) P-BLEU $\downarrow$ (De-En)
(b) distinct-2 $\uparrow$ (De-En)
(c) distinct-3 $\uparrow$ (De-En)
(d) P-BLEU $\downarrow$ (Ru-En)
(e) distinct-2 $\uparrow$ (Ru-En)
(f) distinct-3 $\uparrow$ (Ru-En)
Figure 2: Max BLEU (Oracle score), pairwise-BLEU, distinct-n on WMT’19 De-En and Ru-En. The size of the output $k$ is set to 4. The results of distinct-1 and $k=8,12$ are present in Appendix.
(a) P-BLEU $\downarrow$ (De-En)
(b) P-BLEU $\downarrow$ (De-En)
(c) P-BLEU $\downarrow$ (De-En)
(d) distinct-2 $\uparrow$ (De-En)
(e) distinct-2 $\uparrow$ (De-En)
(f) distinct-2 $\uparrow$ (De-En)
Figure 3: Evaluation of the DMBR and KMBR. Mean BLEU, pairwise-BLEU (P-BLEU), distinct-2 on WMT’19 De-En are reported. (Left) Comparison with varying numbers of the outputs: $k=4,8,12$. (Middle) Comparison of DMBR, KMBR with varying numbers of samples: $N=32,64,128$. (Right) Comparison of the choice of the sampling algorithms: ancestral sampling, nucleus sampling, top-$k$ sampling, and epsilon sampling.
@@ 4.1 Machine Translation
We set the number of outputs to be $k=4,8$, and $12$. We compare the performance of sampling algorithms, diverse beam search, and the proposed methods. For sampling algorithms, we evaluate ancestral sampling, nucleus sampling ($p=0.9$), top-$k$ sampling ($k=10$), and epsilon sampling ($\epsilon=0.02$) Holtzman et al. (2020); Fan et al. (2018); Hewitt et al. (2022). For DBS, we set the number of groups and the beam width to be $k$. The diverse penalty for DBS is set to $0.0$ (beam search), $0.2,0.5,1.0,2.0$, and $5.0$. For DMBR and KMBR, we sample 128 sentences per source sentence with epsilon sampling ($\epsilon=0.02$) Hewitt et al. (2022); Freitag et al. (2023). The diversity penalty $\lambda$ is set $0.0$ (vanilla MBR), $0.1,0.3,0.5,1.0$, and $2.0$.
DMBR achieves higher diversity
@@
All the evaluation metrics are summarized in Table 7, 8, and 9 in the Appendix. Figure 1 shows the P-BLEU and distinct-2, 3 as a function of mean BLEU score. The results show that DMBR achieves higher diversity (lower P-BLEU and higher distinct-2, 3) than DBS and sampling algorithms with the same mean BLEU score.
@@
DMBR achieves more flexibility than DBS on the quality-diversity trade-off
@@
We observe that DBS does not increase the diversity by increasing the diversity penalty larger than 10.0. On the other hand, the diversity of DMBR continue to increase with larger $\lambda$, achieving higher maximum diversity than DBS. DMBR also achieves higher diversity than the sampling algorithms.
DMBR achieves higher diversity than the underlying sampling algorithm
By explicitly choosing samples to maximize the diversity, DMBR achieves higher diversity than the underlying sampling algorithm for all four sampling algorithms we evaluate (Figure 1, 2(c), and 2(f)). This is contrary to many of the beam search-based diversity decoding algorithms as they often fail to improve the diversity upon random sampling. The results show the flexibility of the DMBR to control the trade-off between quality and diversity.
DMBR achieves higher Oracle score than vanilla MBR
@@
Figure 2 shows the P-BLEU and distinct-2, 3 as a function of max BLEU score (i.e. Oracle score). We observe that DMBR achieves a slightly lower maximum BLEU score, yet with higher diversity. Interestingly, we observe that DMBR and KMBR achieve higher Oracle scores than vanilla MBR. This shows the potential of DMBR to further improve the quality of the output of MBR. This is analogous to DBS achieving a higher Oracle score than beam search Vijayakumar et al. (2016). KMBR achieves the highest maximum BLEU score over all algorithms. Although KMBR achieves a lower mean BLEU score than DMBR, it achieves a maximum BLEU score higher than DBS while keeping higher diversity.
DMBR does not scale to the number of outputs with a fixed amount of samples
@@
Figure 2(a) and 2(d) shows the quality and diversity trade-off with varying output size $(k=4,8,12)$. We observe that with a fixed amount of samples ($N=128$), the diversity decreases with a larger output size whereas DBS scales with a larger output size.
DMBR improves with the number of samples
@@
Figure 2(b) and 2(e) shows the performance of DMBR and KMBR with varying number of samples ($N=32,64,128$). We observe that the quality-diversity trade-off is improved with a larger sample size, yet it does not significantly improve by doubling the sample size. The results indicate that one can improve both quality and diversity by increasing the sample size, but it comes at the cost of inference time. Note that the computational complexity of DMBR is quadratic to the number of samples.
Choice of the underlying sampling algorithm for DMBR matters
@@
Figure 2(c) and 2(f) shows the comparison of DMBR with varying sampling algorithms. We evaluate ancestral sampling, nucleus sampling ($p=0.9$), top-$k$ sampling ($k=10$), and epsilon sampling ($\epsilon=0.02$) Holtzman et al. (2020); Fan et al. (2018); Hewitt et al. (2022). The temperature is set to 1 for all the runs. We observe that the choice of the sampling algorithm changes the Pareto front of the DMBR. Using less biased sampling algorithms such as ancestral sampling, DMBR improves the diversity, and using more focused sampling algorithms such as epsilon sampling, DMBR improves the mean quality of the outputs.
@@
(b) distinct-2 $\uparrow$ (MS COCO)
(c) distinct-3 $\uparrow$ (MS COCO)
(d) P-BLEU $\downarrow$ (SQuAD v2)
@@
(f) distinct-3 $\uparrow$ (SQuAD v2)
(g) P-BLEU $\downarrow$ (CommonGen)
(h) distinct-2 $\uparrow$ (CommonGen)
(i) distinct-3 $\uparrow$ (CommonGen)
(j) P-BLEU $\downarrow$ (XSum)
(k) distinct-2 $\uparrow$ (XSum)
(l) distinct-3 $\uparrow$ (XSum)
Figure 4: Evaluation of the pairwise-BLEU (P-BLEU) and distinct-2, 3 as a function of mean BLEU, METEOR, and ROUGE-L. The number of the outputs $k$ is $4$.

### v2

@@ Abstract
One of the most important challenges in text generation systems is to produce outputs that are not only correct but also diverse. Recently, Minimum Bayes-Risk (MBR) decoding has gained prominence for generating sentences of the highest quality among the decoding algorithms. However, existing algorithms proposed to generate diverse outputs are predominantly based on beam search or random sampling, thus their output quality is capped by these underlying decoding algorithms. In this paper, we investigate an alternative approach – we develop diversity-promoting decoding algorithms by enforcing diversity objectives to MBR decoding. We propose two variants of MBR; (i) Diverse MBR (DMBR) that adds a diversity penalty to the decoding objective and (ii) $k$-medoids MBR (KMBR) that reformulates the decoding task as a clustering problem. We evaluate DMBR and KMBR on a variety of directed text generation tasks using encoder-decoder models and a language model with prompting. The experimental results show that the proposed methods achieve a better trade-off than the diverse beam search and sampling algorithms overall. Our code is available at https://github.com/CyberAgentAILab/diverse-mbr/.
Generating Diverse and High-Quality Texts by
Minimum Bayes Risk Decoding
Yuu Jinnai, Ukyo Honda, Tetsuro Morimura, Peinan Zhang CyberAgent {jinnai_yu,honda_ukyo,morimura_tetsuro,zhang_peinan}@cyberagent.co.jp
@@ 1 Introduction
In this paper, we develop diverse decoding methods by extending Minimum Bayes Risk (MBR) decoding Goel and Byrne (2000); Kumar and Byrne (2002, 2004); Eikema and Aziz (2022). MBR decoding is shown to generate higher quality sentences than random sampling and beam search in directed text generation tasks including machine translation, text summarization, and data-to-text Freitag et al. (2022); Suzgun et al. (2023). The procedure of the MBR decoding consists of the following steps. First, it samples a set of candidate outputs from the probability model. Then, it computes the similarity of each sequence to the others according to a utility function. Finally, it selects the sequence that maximizes the expected utility over the sequences. A naive approach to generate $k$ outputs using MBR is to select the top-$k$ outputs with the highest expected utility. However, it tends to select a set of similar sentences with a large overlap (see Appendix C for examples).
@@ 2 Background
In this paper, we consider a variant of the set decoding problem Meister et al. (2021) where the task is to generate a set of $k$ sentences $H$ that maximizes the sum of $P_{\mathrm{human}}$ and the diversity according to the preference of human $d_{\mathrm{human}}$:
@@
In this paper, we use automated evaluation metrics to approximate $d_{\mathrm{human}}$. In particular, we consider P-BLEU, distinct-n, and P-SentBERT as measures of diversity Shen et al. (2019); Li et al. (2016); Reimers and Gurevych (2019).
@@ 2.1 Decoding Algorithms for Diversity
There have been two major approaches to diversity-aware text decoding: random sampling and diversity-aware beam search.
Random sampling.
Random sampling is commonly used to generate diverse outputs in both directed and open-ended text generation tasks. A simple solution is to use an ancestral sampling with a temperature parameter to control the stochasticity of the sampling. There have been a lot of studies on biasing the ancestral sampling to generate higher quality outputs while maintaining the diversity of the randomized algorithm. Prior work shows that top-$k$ sampling that restricts the sampling to the $k_{\mathrm{top}}$ most likely tokens at each step is a better alternative to controlling the temperature in a story generation task Fan et al. (2018); Ippolito et al. (2019). Nucleus sampling Holtzman et al. (2020) is similar to top-$k$ sampling and has shown to be more effective than top-$k$ sampling in WebText dataset Radford et al. (2019). Nucleus sampling truncates all tokens except those in the nucleus, the smallest possible set of tokens that covers a fraction $p$ of the model probability. Epsilon sampling Hewitt et al. (2022) is also a variant of ancestral sampling that truncates tokens whose probability is less than a threshold. They are shown to be more effective than nucleus sampling in the WebText dataset. However, these random samplings improve the diversity of outputs at the expense of the quality of outputs (Ippolito et al., 2019).
Diversity-aware beam search.
Another line of work is to generate a diverse set of outputs by introducing diversity objectives to the beam search. Beam search is known to produce higher quality sequences than random sampling in a wide range of tasks Graves (2012); Sutskever et al. (2014). Diversity has been induced in the beam search procedure in various forms. Li and Jurafsky (2016) propose to add a diversity constraint to standard beam search so that the number of descendants from the same parent hypothesis is capped at some upper bound. Noisy Parallel Approximate Decoding induces noise to the hidden state of the decoder to generate randomized outputs Cho (2016). Diverse beam search (DBS) adds a diversity term to the reranking objective, penalizing sequences with a small Hamming distance to other groups of the sequences Vijayakumar et al. (2016, 2018). Iterative beam search runs beam search multiple times with a constraint that any partial hypothesis that has been generated in previous iterations is prohibited Kulikov et al. (2019). Clustered beam search prunes similar sequences by clustering the candidates using a word embedding at each decoding step Tam (2020). Post-Decoding Clustering (PDC) encourages diversity by running the clustering after generating a large number of outputs, selecting good candidates from each cluster Kriz et al. (2019); Ippolito et al. (2019). Best-$k$ search maintains best-first queue in a course of beam search, resulting in higher quality and diversity outputs Xu et al. (2023a).
@@ 2.2 Minimum Bayes Risk (MBR) Decoding
One of the most common decision rules to solve the decoding problem is maximum-a-posteriori (MAP) decoding such as beam search. MAP decoding finds the most probable translation under the model.
@@
$\mathbf{h}^{\mathrm{MBR}}=\operatorname*{arg\,max}_{\mathbf{h}\in\mathcal{H}_{\mathrm{cand}}}\frac{1}{N}\sum_{\mathbf{y}\in\mathcal{H}_{\mathrm{ref}}}u(\mathbf{h},\mathbf{y}),$ (7)
where $N=|\mathcal{H}_{\mathrm{ref}}|$. Standard practice is to use the same set of hypotheses for the candidate pool ($\mathcal{H}$) and the reference pool ($\mathcal{H}_{\mathrm{ref}}$), therefore $\mathcal{H}=\mathcal{H}_{\mathrm{ref}}$.
@@ 3 Minimum Bayes Risk Decoding with Diversity
However, it results in a set of similar sentences with a large overlap. In fact, it often finds almost duplicated sentences in machine translation tasks (Tables 2 and 3 in Appendix C).
@@ 3.1 Diverse MBR (DMBR)
$H^{\mathrm{DMBR}}=$
$\;\;\operatorname*{arg\,max}_{\stackrel{{\scriptstyle H\subseteq\mathcal{H}_{\mathrm{cand}}}}{{|H|=k}}}\sum_{\mathbf{h}\in H}\left(\frac{1}{N}\sum_{\mathbf{y}\in\mathcal{H}_{\mathrm{ref}}}u(\mathbf{h},\mathbf{y})\right)+d(H)$ (9)
@@
$d(H)$ is the diversity objective that penalizes according to a user-defined diversity objective. In this paper, we aim to promote diversity by minimizing the pairwise similarity among the outputs. Since the utility function typically computes the similarity between each pair of texts, we minimize the following pairwise utility:
@@ 3.2 $k$-Medoids MBR (KMBR)
As an alternative approach to promote diversity, we propose $k$-Medoids MBR (KMBR) decoding. $k$-Medoids is a clustering problem similar to $k$-means in that it chooses centers to minimize the total distance of data points to the closest center points. The difference is that $k$-Medoids needs to choose actual data points as centers Rdusseeun and Kaufman (1987); Kaufman and Rousseeuw (2009). Intuitively, $k$-Medoids center points are supposed to be representative of different clusters of hypotheses. Thus, picking the center points for clusters is likely to result in a set of diverse and high-quality hypotheses. KMBR can be understood as a generalization of the vanilla MBR decoding, which is solving the $1$-Medoid problem to find the single most centered hypothesis out of the sampled hypotheses Jinnai and Ariu (2024). We use the negative utility as a distance and consider the problem of picking a set of $k$ center points. The total distance from the reference set (= data points) to the picked center points is minimized as follows:
$H^{\mathrm{KMBR}}$ $=\operatorname*{arg\,max}_{\stackrel{{\scriptstyle H\subseteq\mathcal{H}_{\mathrm{cand}}}}{{|H|=k}}}\sum_{\mathbf{y}\in\mathcal{H}_{\mathrm{ref}}}\min_{\mathbf{h}\in H}-u(\mathbf{h},\mathbf{y}).$ (11)
$k$-medoids can be used with arbitrary dissimilarity measures. Because the $k$-medoids problem is NP-hard to solve exactly, we deploy Partition Around Medoids (PAM) to compute $H^{\mathrm{KMBR}}$ approximately Park and Jun (2009). Similarly to DMBR, KMBR needs to compute the utility function for every pair of hypotheses so it is still slower than MBR even with the approximation algorithm.
@@
(b) P-BLEU $\downarrow$ (Ru-En)
(c) distinct-2 $\uparrow$ (De-En)
(d) distinct-2 $\uparrow$ (Ru-En)
Figure 1: Evaluation of P-BLEU and distinct-2 as a function of mean BLEU on WMT’19 De-En and Ru-En. The size of the outputs $k$ is set to 4. $\uparrow$ and $\downarrow$ denote that larger and smaller are better in diversity, respectively.
(e) P-BLEU $\downarrow$ (De-En)
(f) distinct-2 $\uparrow$ (De-En)
Figure 2: Evaluation of P-BLEU, distinct-2 as a function of max BLEU (Oracle score) on WMT’19 De-En. The size of the outputs $k$ is set to 4.
@@ 4 Experiments
We evaluate DMBR and KMBR in machine translation, image captioning, question generation, generative common sense reasoning, and text summarization. All experiments use BERTScore Zhang* et al. (2020) as the utility function of the MBR. We compare the performance of the sampling algorithms, beam search, diverse beam search (DBS) Vijayakumar et al. (2018), MBR, diverse MBR (DMBR), and $k$-Medoids MBR (KMBR).
We evaluate the quality and the diversity of the generated texts. As a quality metric, we report the mean, max, and min quality over each set of $k$ sentences measured by BLEU, ROUGE-L, or METEOR Papineni et al. (2002); Lin and Och (2004); Banerjee and Lavie (2005). We use distinct-n Li et al. (2016) and pairwise-BLEU (P-BLEU) Shen et al. (2019) as diversity metrics.
Due to limitations in computational resources, we run experiments using the first 1000 entries of the dataset for all the experiments in this paper. We use Huggingface’s Transformers library for running all the experiments Wolf et al. (2020). We initialize the clusters for PAM used in KMBR by k-medoids++ with maximum iterations set to $300$. For reproducibility, all the experiments are conducted using publicly available pretrained models and datasets. We use sacreBLEU system Post (2018) to compute BLEU scores.
Examples of generations are shown in Appendix C. See Appendix E for additional figures and the summary of the experimental results in tables.
(a) P-BLEU $\downarrow$ with varying #outputs
(b) distinct-2 $\uparrow$ with varying #outputs
Figure 3: Evaluation of DMBR and KMBR with varying number of outputs ($k\in\{4,8,12\}$). Mean BLEU, P-BLEU, and distinct-2 on WMT’19 De-En are reported.
(a) P-BLEU $\downarrow$ with varying #samples
(b) distinct-2 $\uparrow$ with varying #samples
Figure 4: Evaluation of DMBR and KMBR with varying number of samples ($N\in\{32,64,128\}$). Mean BLEU, P-BLEU, and distinct-2 on WMT’19 De-En are reported.
(a) P-BLEU $\downarrow$ with varying sampling
(b) distinct-2 $\uparrow$ with varying sampling
Figure 5: Evaluation of DMBR and KMBR using varying sampling algorithms: ancestral sampling, nucleus sampling, top-$k$ sampling, and epsilon sampling.. Mean BLEU, P-BLEU, and distinct-2 on WMT’19 De-En are reported.
@@ 4.1 Machine Translation
[…]
Due to space limitations, we discuss the essential results here and show the detailed results in Tables 9, 10, and 11 in Appendix E.
DMBR achieves higher diversity than baselines.
@@
Figure 2 shows the P-BLEU and distinct-2 as a function of mean BLEU score. The results show that DMBR achieves higher diversity (lower P-BLEU and higher distinct-2) than DBS and sampling algorithms with the same mean BLEU score.
@@
DMBR achieves more flexibility than DBS on the quality-diversity trade-off.
@@
We observe that DBS does not increase the diversity by increasing the diversity penalty larger than $10.0$ (Table 9 in Appendix). On the other hand, the diversity of DMBR continue to increase with larger $\lambda$, achieving higher maximum diversity than DBS. DMBR also achieves higher diversity than the sampling algorithms.
DMBR achieves higher oracle score than vanilla MBR.
@@
The Oracle score indicates the score of the highest-scoring output in a set of $k$ outputs Vijayakumar et al. (2018). It is intended to evaluate how well the obtained diversity benefits in producing output close to a particular correct answer. Figure 2 shows the P-BLEU and distinct-2 as a function of max BLEU score (i.e., Oracle score). We observe that DMBR achieves a slightly lower max BLEU score than DBS, yet with higher diversity. Interestingly, we observe that DMBR and KMBR achieve higher max scores than vanilla MBR (Table 9). This shows the potential of DMBR to further improve the quality of the output of MBR. This is analogous to DBS achieving a higher Oracle score than beam search Vijayakumar et al. (2016). KMBR achieves the highest max BLEU score over all algorithms compared.
DMBR outperforms DBS with varying output sizes.
@@
Figure 3 shows the quality and diversity trade-off with varying output size $k\in\{4,8,12\}$. DBS shows no degradation of P-BLEU but shows a lower distinct-2 score with a larger output size. Overall, we observe that DMBR outperforms DBS in both diversity metrics with all the output sizes we evaluated.
DMBR improves with a larger number of samples.
@@
Figure 4 shows the performance of DMBR and KMBR with varying numbers of samples $N\in\{32,64,128\}$. We observe that the quality-diversity trade-off is improved with a larger sample size, yet it does not significantly improve by doubling the sample size. The results indicate that one can improve both quality and diversity by increasing the sample size, but it comes at the cost of inference time. Note that the computational complexity of DMBR is quadratic to the number of samples.
Choice of the underlying sampling algorithm matters.
@@
Figure 5 shows the comparison of DMBR with varying sampling algorithms. We evaluate ancestral sampling, nucleus sampling with $p\in\{0.8,0.9,0.95\}$, top-$k$ sampling with $k_{\mathrm{top}}\in\{5,10,50\}$, and epsilon sampling with $\epsilon\in\{0.01,0.02,0.04\}$ Holtzman et al. (2020); Fan et al. (2018); Hewitt et al. (2022). The temperature is set to 1 for all the runs. We observe that the choice of the sampling algorithm changes the Pareto front of the DMBR. Using less biased sampling algorithms such as ancestral sampling, DMBR improves the diversity, and using more focused sampling algorithms such as epsilon sampling, DMBR improves the mean quality of the outputs.
@@
(b) P-BLEU $\downarrow$ (SQuADv2)
(c) P-BLEU $\downarrow$ (CommonGen)
(d) distinct-2 $\uparrow$ (MS COCO)
@@
(f) distinct-2 $\uparrow$ (CommonGen)
(g) P-SentBERT $\downarrow$ (MS COCO)
(h) P-SentBERT $\downarrow$ (SQuADv2)
(i) P-SentBERT $\downarrow$ (CommonGen)
Figure 6: Evaluation of P-BLEU, distinct-2, and P-SentBERT as a function of mean BLEU (MS COCO) and METEOR (SQuADv2 and CommonGen). The number of the outputs $k$ is $4$.